REPOST: Oil and gas. An outboard motor calls for a fuel mixture that has a gas-to-oil ratio of 50 to 1. How many pints of oil should be added to 6 gallons of gasoline?
Answer by stanbon(48558)

An outboard motor calls for a fuel mixture that has a gas-to-oil ratio of 50 to 1. How many pints of oil should be added to 6 gallons of gasoline?
1 gal. = 4 quarts
1quart = 2 pints
So, 1 gallon = 8 pints
6 gallons = 48 pints.
gas/oil = 50/1 = 48 pints gas/x pints of oil
EQUATION:
x/48= 1/50
50x=48
x= 0.96 pint of oil
Cheers,
Stan H.